Suppose that you are considering the acquisition of a factory that specializes in the manufacture of...

  1. Suppose that you are considering the acquisition of a factory that specializes in the manufacture of nose hair trimmers. The Trimmer is marketed on television for $9.99, is battery operated, and uses a swirling blade much like the Norelco razor. The factory manufactures about 1,000,000 units per year, sells the trimmer to the wholesaler for $6.37 per unit, and has a cost of manufacture of $1.87 per unit. If the factory must rent space for $1,000,000 per year, the expected life of the equipment is 15 years, demand is estimated to be steady, and the tax rate is 28%, what is the NPV if you pay $25,000,000 for the factory and your required return is 15% (after-tax)? Show a detailed pro-forma financial model. You should not need Excel to do this problem.
Answer #1
Sales 6,370,000
COGS 1,870,000
SGA 1,000,000
EBT 3,500,000
Tax (28%)      980,000
Profits 2,520,000

Calculate the annual profits from the factory as shown above.

Now. present value of profits for 15 years can be calculated using PV function on a calculator

N = 15, PMT = 2,520,000, I/Y = 15%, FV = 0,

=> Compute PV = $14,735,372.65

NPV = 14,735,372.65 - 25,000,000 = -10,264,627.35